OpenStudy (anonymous):

A 20 N bucket is to be raised at a constant speed of 75 cm/s by a rope. How many kilograms of cement can be put into this bucket without breaking the rope if it can withstand a maximum tension of 445 N?

OpenStudy (anonymous):

if the bucket is to raised at constant speed the force on the tension must be equal to the weight

OpenStudy (anonymous):

does it make sense to u?

OpenStudy (anonymous):

therefore 445=20+mg where g=9.8 u can now solve for ur mas

OpenStudy (anonymous):

@carbone r u there???????

OpenStudy (anonymous):

445 = 20 + mg 425 = mg taking g as 10 m = 42.5Kg

OpenStudy (anonymous):

@ yahoo that's it ; @ aperogalics it that the right answer?

OpenStudy (anonymous):

@hubertH it is right answer:)

OpenStudy (anonymous):

cool guys